OpenStudy (anonymous):

Find the amplitude of y = -4 sin x.

OpenStudy (raffle_snaffle):

Is this precalc?

OpenStudy (raffle_snaffle):

The amplitude is -4

OpenStudy (anonymous):

wouldn't it just be 4?

OpenStudy (raffle_snaffle):

^^^Yes you are correct it's just 4.

OpenStudy (anonymous):

ahahahahahah, yay :)

terenzreignz (terenzreignz):

Okay, let's have a look at your generic sine wave (note that it makes no difference if it were sin instead of cos, the basic principles remain the same)

terenzreignz (terenzreignz):

\[\Large y = \color{blue}a \cos(\color{orange}px + \color{green}q)+ b\]

terenzreignz (terenzreignz):

The amplitude is given by \[\Large |\color{blue}a|\]

terenzreignz (terenzreignz):

The amplitude is the highest value your function could possibly take (without the shift) And since cosine (or sine) can only ever be as high as 1, then, the entire function can only ever be as high as you coefficient of your trig-function.

terenzreignz (terenzreignz):

If you're ever asked what the PERIOD is, then consider that coefficient of the x (inside the cos or sin) Divide \(\large 2\pi\) by that coefficient and you'll have the period. \[\Large T = \frac{2\pi }p\]